# Modelling reliability of reversible circuits with 2D second-order cellular automata

### 20 Dec 2023 v1

@@ 4 Discussion and conclusions
Despite of impossibility to use $E_{1,3}^{I\!I}$ for construction of reversible circuits this CA is included in the comparison due to known remarkable properties such as linearity and self-replication of any pattern in $E_{1,3}$ [25]. All CA used here for construction of second-order RCA could be informally treated as some non-linear modifications of the Fredkin replicator $E_{1,3}$.
@@
The analytical formulas Eq. (11) obtained due to linearity of $E_{1,3}^{I\!I}$ also can be extended to $C_{0}E_{1}^{\,I\!I}$ [31] (and $C_{0}E_{1,3}^{\,I\!I}$) to analyse simple example of damage distribution. Oscillating behavior of formal dimension ${\mathfrak{D}}$ understandable from such expressions provides some hints about possible issues with instability in damage distribution for considered family of RCA partially confirmed by numerical experiments.
@@
Due to such instability the RCA family has high sensitivity to errors, because even local defects or improper functionality of single gate finally leads to failure of the entire circuit. However, such behavior also provides simple error indication. A fault almost inevitably produces quickly recorded growth in total number of non-empty cells signaling a failure and preventing the completion of calculations with an incorrect result.
@@
Similarly with BBM (billiard ball model of computations) RCA discussed in this work do not use any ‘wires’ for motion between gates. However, similar models with ‘wires’ also can be designed using generalization of second-order RCA [47] implemented in software mentioned earlier [32]. Such RCA have five states with additional one for construction of ‘wires’. That could reduce sensitivity to external noise, because all cells around wires are passive and may not participate in damage distribution. Reliability analysis for such RCA with damage distribution only along wires should be discussed elsewhere.
@@
The existence of alternative models with ‘wires’ also justifies earlier informal name ‘wireless ants’ for part of RCA family discussed here and abbreviations for local rules such as WlAnt, WlAnt-2, WlAnt-3 and WlAnt-4 in software [32] for $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ respectively. The local rules for $C_{0}E_{1}^{\,I\!I}$ (WlSnake) and $C_{0}E_{1,3}^{\,I\!I}$ (WlSnake-3) are implemented in the same software [32]. All rules are also converted to so-called ‘tree’ format used for CA simulations in more common software [33].

### 19 Mar 2024 v2

@@ 4 Discussion and conclusions
Despite of impossibility to use $E_{1,3}^{I\!I}$ for construction of reversible circuits this CA is included in the comparison due to known remarkable properties such as linearity and self-replication of any pattern in $E_{1,3}$ Gard . All CA used here for construction of second-order RCA could be informally treated as some non-linear modifications of the Fredkin replicator $E_{1,3}$.
@@
The analytical formulas Eq. (11) obtained due to linearity of $E_{1,3}^{I\!I}$ also can be extended to $C_{0}E_{1}^{\,I\!I}$ AV14 (and $C_{0}E_{1,3}^{\,I\!I}$) to analyse simple example of damage distribution. Oscillating behaviour of formal dimension ${\mathfrak{D}}$ understandable from such expressions provides some hints about possible issues with instability in damage distribution for considered family of RCA partially confirmed by numerical experiments.
@@
Due to such instability the RCA family has high sensitivity to errors, because even local defects or improper functionality of single gate finally leads to failure of the entire circuit. However, such behaviour also provides simple error indication. A fault almost inevitably produces quickly recorded growth in total number of non-empty cells signaling a failure and preventing the completion of calculations with an incorrect result.
On the other hand, the serious damage from even a single error, along with the difficulty of isolation other elements from its influence, lead to problems with longer reliable work of reversible circuits with many gates partially consistent with conclusion drawn from altentative model with probabilistic RCA RCAnoise .
The problem with sensitivity to errors may be partially related with constructions of considered RCA. They are adapted for simpler design of logic gates, but not for reliability. The local rules of RCA considered in this work provide a natural way to create small mobile configurations that easily change direction of movement. At the same time, close similarity with RCA $E_{1,3}^{I\!I}$ derived from Fredkin replicator could result extra instability due to variance of formal dimension ${\mathfrak{D}}$ already discussed earlier in Sec. 3.
Only few RCA from rather big families such as CEOT are discussed. However for construction of reversible logic circuits even one of them such as $C_{0}E_{1,2}^{\,I\!I}$ would be enough to demonstrate examples with both conserved and variable number of signals. Other RCA discussed here are useful for analysis of reliability and damage distribution.
For all RCA discussed here computational experiments demonstrate too serious problems for whole circuits even from single errors. Possibility to reduce such effects could be considered as a reasonable research direction.
Perhaps, the choice of local rules aimed at greater stability of configurations could reduce sensitivity to errors. For example, even small difference between $C_{0}E_{1,2}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ hinders the damage spread after side impact with zero delay, cf Figure 3 with Figure 5 for $\Delta=0$ in A.
The $C_{0}E_{1,2o}^{\,I\!I}$ belongs to even wider family, than CEOT. It is second-order RCA derived from isotropic inner-independent CA. The total number of isotropic CA mentioned earlier is $2^{102}\simeq 5\times 10^{30}$ and number of inner-independent isotropic CA is $2^{51}\simeq 2\times 10^{15}$. Thus, brute force method is hardly appropriate to search for more reliable RCA.
There is simple measure of complexity for such local rules from CET and CEOT family based on counting number $N_{p}$ of mentioned in Sec. 2.2 pairs $p=(n_{c},n_{e})$ with $f(p)=1$ producing change of central cell. For example, $N_{p}=1$ for $C_{0}E_{1}$ or $C_{0}E_{1}^{\,I\!I}$ and $N_{p}=2$ for $C_{0}E_{1,2}^{\,I\!I}$. For isotropic rule similar number $N_{I}$ could be used, but instead number of different configurations up to rotations and reflections should be used.
For example pair $(c_{c}=0,n_{e}=2)$ includes two different isotropic configurations with neighbouring cells either opposite or with touching corners. The first case is used in definition of $C_{0}E_{1,2o}^{\,I\!I}$. Thus, $N_{I}=1$ for $C_{0}E_{1}^{\,I\!I}$, $N_{I}=2$ for $C_{0}E_{1,2o}^{\,I\!I}$ and $N_{I}=3$ for $C_{0}E_{1,2}^{\,I\!I}$.
For such complexity measures the universal computation can be obtained already for ‘simplest’ RCA $C_{0}E_{1}^{\,I\!I}$, $C_{0}E_{1,2o}^{\,I\!I}$ with $N_{I}=1,2$. More complex local rules may be necessary for investigation of stability and damage distribution, but each new pair or configuration included in the rule may produce destruction of initially stable static elements used for signal routing or undesirable interactions of signals with such elements. Such problems with more complex local rules could be partially avoided by using special passive ‘mirrors’ or active ‘wires’ for signal routing.
@@
Similarly with BBM (billiard ball model of computations) RCA discussed in this work do not use any ‘wires’ for motion between gates. Analogue models with ‘wires’ could be designed using generalization of second-order RCA AV13 implemented in software mentioned earlier CART . Such RCA have five states with additional one for construction of ‘wires’. That could reduce sensitivity to external noise, because all cells around wires are passive and may not participate in damage distribution. Reliability analysis for such RCA with damage distribution only along wires should be discussed elsewhere.
@@
The existence of alternative models with ‘wires’ also justifies earlier informal name ‘wireless ants’ for some RCA discussed here and abbreviations for local rules in software CART such as WlAnt, WlAnt-2, WlAnt-3 and WlAnt-4 for $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ respectively. The local rules for $C_{0}E_{1}^{\,I\!I}$ (WlSnake) and $C_{0}E_{1,3}^{\,I\!I}$ (WlSnake-3) are implemented in the same package. All rules are also converted to so-called ‘tree’ format used for CA simulations in more common software GoL .
Appendix A Modelling of signals collision
Collisions of signals with different delays $\Delta$ are shown for discussed in this work RCA on Figures 15 for $t=0$, $t=7$ and $t=14$. The configurations for $t=0$ and $t=7$ are equal for all RCA and shown only on Figure 1. Due to symmetry of picture with respect to reflections about the diagonal only six values $0\leq|\Delta|\leq 5$ are shown for each case. The value $|\Delta|=5$ is the minimal one without any interaction.
Figure 1: Collisions with different delays in $C_{0}E_{1}^{\,I\!I}$
Figure 2: Collisions with different delays in $C_{0}E_{1,3}^{\,I\!I}$
Figure 3: Collisions with different delays in $C_{0}E_{1,2}^{\,I\!I}$
Figure 5: Collisions with different delays in $C_{0}E_{1,2o}^{\,I\!I}$
For RCA $C_{0}E_{1}^{\,I\!I}$ and $C_{0}E_{1,3}^{\,I\!I}$ of the first type sometime referred earlier as ‘snakes’ depicted on Figure 1 and Figure 2 the value $\Delta=0$ corresponds to reflection of signals followed by motion in opposite direction. The values $1\leq|\Delta|\leq 3$ produce a signal expanding in two opposite directions together with initial one. The value $|\Delta|=4$ seems produce more critical damage with front expanding in all possible directions.
The same value $|\Delta|=4$ also looks critical for second type of RCA $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ sometimes referred earlier as ‘ants’ depicted on Figure 3 – Figure 5, but only for $C_{0}E_{1,2o}^{\,I\!I}$ value $\Delta=0$ produces simple reflection and for other four RCA of such type $\Delta=0$ causes more critical omnidirectional damage spread. The values $1\leq|\Delta|\leq 3$ for all four RCA of such type due to split of one signal after collision produce three signals moving apart with apparently smaller damage.
Both $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{+}^{\,I\!I}$ produce the same damage patterns Figure 4 with ${\mathfrak{D}}_{\infty}=2$ for $\Delta=0$ or ${\mathfrak{D}}_{\infty}=1$ for $|\Delta|=4$, despite of different behaviour for initial configuration with single red cell discussed earlier, see Figure 7 and Figure 8.
Appendix B Modelling of signals interaction
An elementary interaction of signals is depicted on Figures 15. The configurations for $t=0$ and $t=7$ again are equal for all RCA and shown only on Figure 1. The main difference in comparison with Figures 15 is $2\times 2$ static blue block (cells with state 3) for proper signal motion after collision. The interaction between signals again corresponds to $-4\leq\Delta\leq 4$ and six values $0\leq|\Delta|\leq 5$ are enough for each case due to symmetry with respect to reflections about the diagonal.
Figure 1: Interactions with different delays in $C_{0}E_{1}^{\,I\!I}$
Figure 2: Interactions with different delays in $C_{0}E_{1,3}^{\,I\!I}$
Figure 3: Interactions with different delays in $C_{0}E_{1,2}^{\,I\!I}$
Figure 5: Interactions with different delays in $C_{0}E_{1,2o}^{\,I\!I}$
For $1\leq|\Delta|\leq 3$ after interaction two signals move along parallel paths with relative delay $|\Delta|-1$. For elements similar with switch gate Figure 2 proper value is $|\Delta|=1$ taking into account necessity for repeated interaction. The value $|\Delta|=4$ produces omnidirectional damage front for all RCA discussed in the paper and $\Delta=0$ is related with similar problem for $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{+}^{\,I\!I}$. For $C_{0}E_{1}^{\,I\!I}$, $C_{0}E_{1,3}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ the same value $|\Delta|=0$ causes reflection of both signals after interaction.
Both $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{+}^{\,I\!I}$ produce the same damage fronts with ${\mathfrak{D}}_{\infty}=1$, see Figure 4. Despite of some resemblance for $\Delta=0$ of initial time steps shown on Figure 3, ${\mathfrak{D}}_{\infty}=2$ for $C_{0}E_{1,2}^{\,I\!I}$ not only for $|\Delta|=4$, but also for $\Delta=0$ with shape of damage described as diamond-like area with rectangular empty space due to initial impact of $2\times 2$ static blue block.